OpenStudy (anonymous):

2 integrate sqare root of [9-2x }, limits 0 to 1

OpenStudy (tkhunny):

This? \(\int\limits_{0}^{1}\sqrt{9-2x}\;dx\)

OpenStudy (tkhunny):

You can substitute u = 9 - 2x or, if your are hardy enough, just do it cold and remember your chain rule.